OpenStudy (anonymous):

An obtuse triangle with area 12 has two sides of lengths 4 and 10. Find the length of the third side. (There are two answers.) (Use law of cosines)

OpenStudy (jamesj):

ok. What's the law of cosines?

OpenStudy (anonymous):

c^2 = a^2 + b^2 - 2abCosC

OpenStudy (jamesj):

right. and what's a formula for the area of the triangle?

OpenStudy (jamesj):

...preferably a formula using the variables in the law of cosines formula you just wrote down

OpenStudy (anonymous):

(1/2)(B)(H)

OpenStudy (jamesj):

We want a formula using a,b,c and C.

OpenStudy (jamesj):

Draw yourself a diagram.

OpenStudy (anonymous):

idk, that's y I need help

OpenStudy (jamesj):

one sec. making one

OpenStudy (anonymous):

I don't understand how u got 26cosC

OpenStudy (jamesj):

2b. not 6, b, the letter b.

OpenStudy (anonymous):

James - It is sinG = h/4, right?

OpenStudy (jamesj):

let me fix that.

OpenStudy (anonymous):

I found the angle C to be 53.13 degrees

OpenStudy (anonymous):

?

OpenStudy (jamesj):

Now, use this and the fact that Area = bh/2 to find an expression for sin C, and from that find an expression for cos C. Do not solve for C; there's no need to. Once you have an expression for cos C, you can then use the law of cosines to find little c, the side length.

OpenStudy (anonymous):

I'm confused

OpenStudy (jamesj):

Ok. Agree with my new diagram?

OpenStudy (jamesj):

talk to me or I'm out of here

OpenStudy (anonymous):

I'm here

OpenStudy (anonymous):

yea, I get that part

OpenStudy (anonymous):

what's next

OpenStudy (jamesj):

Ok, now Area = bh/2. Use that relationship to find the value of sin C. From that, find the value of cos C. And then you can use the law of cosines to find little c.

OpenStudy (anonymous):

b in bh/2 is 10, right?

OpenStudy (anonymous):

?

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

I didn't get the answer. I got C = 36.9 degrees. What's next??

OpenStudy (jamesj):

No, stop, wait. What expression do you get for sin C?

OpenStudy (anonymous):

ok, I got 12= (1/2)(10)(4sinC), 12 = 20sinC, sinC = 12/20

OpenStudy (jamesj):

yes, sin C = 3/5. Hence what is cos C. Notice it has two values

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

?

OpenStudy (jamesj):

sin^2 C + cos^2 C = 1. Hence cos C has two values: \[ \cos C = \pm \sqrt{1-\sin^2 C} \]

OpenStudy (anonymous):

so 4/5 = +/- sqrt of 1-sin^2C???

OpenStudy (jamesj):

\[ \cos C = \pm \frac{4}{5} \]

OpenStudy (jamesj):

Now use the law of cosines to find little c.

OpenStudy (anonymous):

\[(4^2) +(10^2)- (2*4*10)(4/5) = c^2 \]

OpenStudy (anonymous):

it didn't work, doesn't match the answer key

OpenStudy (anonymous):

I know, but isn't that how u find the value for one c?

OpenStudy (jamesj):

Nooo.... cos C has two values. And in any case, you need to take the square root to find c.

OpenStudy (jamesj):

yes. Did you take the square root?

OpenStudy (anonymous):

show me how u do it then

OpenStudy (jamesj):

No ... you nut it out. I've already done 90% of this problem for you. You finish it and if it takes you 20 minutes, that's fine. That's how you learn. Figure it out. That's how all of us have to learn. It's painful. But you'll feel better with yourself if you finish it properly.

OpenStudy (anonymous):

I don't get how to do it, u told me to use the law of cosines, I can't find the answer

OpenStudy (anonymous):

nvm, I got it, thanks!

OpenStudy (jamesj):

see? :-) Now, do yourself a favor. Take a blank piece of paper. Solve the whole problem again. When you can do that without looking at this solution or any other version of the solution, then you understand the question. This is what I used to do a lot of High School, and it helped me really learn the material